36 SSM The volume of a cone of height h and base radius r is V-Tr2h/3. A conical vessel of height 25 cm rest- ing on its base of radius 15 cm is filled with water. (a) Find the volume and weight of the water in the vessel. (b) Find the force exerted by the water on the base of the vessel. Explain how this force can be greater than the weight of the water.
Answer #1

Here ,

V = pi * r^2 * h/3

a) volume of the water = pi * 15^2 * 25/3

volume of the water = 5889 cm^3

the volume of the water is 5889 cm^3

weight of water = V * density * g

weight of water = 5889 * 1/1000 * 9.8

weight of water = 57.7 N

b)

force on the base of vessel = pressure * area

force on the base of vessel = 1000 * 9.8 * 0.25 * pi * 0.15^2

force on the base of vessel = 173 N

the force on the base of the vessel is 173 N

this is greater than weight of water as the walls of the container also applies a force downwards on the water.